OpenStudy (anonymous):

HELP! Medal and I WILL become a fan! :) Using complete sentences, explain how you would factor completely x9 –27.

OpenStudy (anonymous):

x^9 - 27 You can start to factor by taking the cubed root of the term cuberoot(x^9 - 27) = x - 3 Are you with me so far?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The reason this is true is because it takes 3 sets of (x-3) to make x^9 - 27 This is because (x-3)(x-3)(x-3) = x^9 - 27 Is that a full enough explanation?

OpenStudy (anonymous):

Yes, thank you so much! :)